OpenStudy (anonymous):

Find the 8th term of the arithmetic sequence: -3.2, -3.4, -3.6, -3.8, ...

OpenStudy (maheshmeghwal9):

\[\huge{\color{green}{T_n=a+(n-1)d.}}\]

Parth (parthkohli):

Notice that the common difference is -3.4 - (-3.2).

OpenStudy (anonymous):

what' di i plug in for d?

OpenStudy (anonymous):

*do

OpenStudy (anonymous):

The eighth term will be when n = 8 an = 91 + (n - 1)(-9) when n = 8 a8 = 91 + (8 - 1)(-9) a8 = 91 + 7(-9) a8 = 91 - 63 a8 = 28 The eighth term is 28

Parth (parthkohli):

Simplify \(-3.4 - (-3.2)\)

Parth (parthkohli):

Careful - @best.shakir is not correct.

OpenStudy (anonymous):

sorry

mathslover (mathslover):

We calculate : common difference : a_2 - _1 ..

mathslover (mathslover):

common diff. = a_2 - a_1

OpenStudy (anonymous):

yes, i just found that out. it's alright. THANK YOU @maheshmeghwal9! The answer is -4.6 :)

Parth (parthkohli):

Yep!

mathslover (mathslover):

good ..

OpenStudy (anonymous):

=D thank you guys!

OpenStudy (maheshmeghwal9):

yw @yodawgiheardyolikedawgs :D